# Reduced Network Cumulative Constraint Violation for Distributed Bandit Convex Optimization under Slater’s Condition

### v1

@@ III-B Performance Analysis for Convex Case
where $c\in({0,1})$, ${\gamma_{0}}\in\big{(}{0,\sigma/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ are constants. Then, for any $T\in{\mathbb{N}_{+}}$,
@@ III-C Performance Analysis for Strongly Convex Case
where ${\gamma_{0}}\in\big{(}{0,\sigma/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ is a constant. Then, for any $T\in{\mathbb{N}_{+}}$,
@@ A. Useful Lemmas
Some preliminary results are given in this subsection.
We first provide some results on the regularized Bregman projection in the following lemma.
@@
In addition, If $\Phi(y)={\|{b-y}\|^{2}}+\langle{a,y}\rangle$, then we know $\Phi$ is a strongly convex function with convexity parameter $\sigma>0$ and ${x_{c}}=\mathop{\arg\min}\limits_{y\in\mathbb{K}}\Phi(y)$. We can obtain
$\|{{x_{c}}-b}\|\leq\frac{{\|a\|}}{\sigma}.$ (30)
@@
We know (29) holds from Lemma 3 in [35]. From the optimality condition, we have
${\|{b-{x_{c}}}\|^{2}}+\langle{a,{x_{c}}}\rangle\leq{\|{b-b}\|^{2}}+2\langle{a,b}\rangle-\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}.$
From ${\|{b-{x_{c}}}\|^{2}}\geq\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}$, and rearranging the above inequality give
$\sigma{\|{b-{x_{c}}}\|^{2}}\leq{\|{b-{x_{c}}}\|^{2}}+\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}\leq\langle{a,b-{x_{c}}}\rangle.$
@@
$\sigma{\|{b-{x_{c}}}\|^{2}}\leq\|a\|\|{b-{x_{c}}}\|.$
From the above inequality and $\sigma>0$, we can obtain (30). ∎
@@
From the Cauchy-Schwarz inequality and (8b), we have
@@
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]\leq\frac{1}{\sigma}\big{(}{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}.$ (47)
@@
and $\sigma>0$ is the uniformly bounded of convexity parameter of the function $\Phi$ as shown in Lemma 1 when $b={x_{i,t}}$ and $a={\alpha_{t}}{{\hat{\omega}}_{i,t+1}}$ for any $i\in[n]$ and $t\in{\mathbb{N}_{+}}$.
@@
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]$ $={\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{{z_{i,t+1}}-{x_{i,t}}}\|]\leq\frac{{{\alpha_{t}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{{{\hat{\omega}}_{i,t+1}}}\|]}}{\sigma}$
$=\frac{{{\alpha_{t}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\hat{\partial}{f_{i,t}}({{x_{i,t}}})+\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}}}\|]}}{\sigma}$
$\leq\frac{{{\alpha_{t}}}}{\sigma}\big{(}{{G_{1}}+{G_{2}}{\gamma_{t}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}$
$=\frac{1}{\sigma}\big{(}{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}.$ (64)
where the second equality holds due to (15b); the last inequality holds due to (8a), (8b) and (51); the last equality holds due to ${\gamma_{t}}{\alpha_{t}}={\gamma_{0}}$. It follows form (64) that (47) holds. ∎
@@
${n_{T}}=\sum\limits_{t=1}^{T}{{\tilde{\varepsilon}_{2}^{2}{{{{G_{1}^{2}}}}}{\alpha_{t}}}}+{m_{T}},{\varepsilon_{2}}=\frac{{4\max\{{1,G_{2}^{2}{{\tilde{\varepsilon}}_{4}}}\}}}{{\min\{{1,\frac{\sigma}{{2{\gamma_{0}}}}}\}}},$
${\varepsilon_{3}}=\frac{{n{{G_{1}}}{G_{2}}{{\tilde{\varepsilon}}_{2}}}}{\sigma},{\varepsilon_{4}}=\frac{{G_{2}^{2}{{\tilde{\varepsilon}}_{2}}{\gamma_{0}}+\sigma}}{\sigma}.$
@@
$\leq n{\tilde{\varepsilon}_{1}}+\frac{{{{\tilde{\varepsilon}}_{2}}}}{\sigma}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\big{(}{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}}}.$ (76)

### v2

@@ III-B Performance Analysis for Convex Case
where $c\in({0,1})$, ${\gamma_{0}}\in\big{(}{0,1/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ are constants. Then, for any $T\in{\mathbb{N}_{+}}$,
@@ III-C Performance Analysis for Strongly Convex Case
where ${\gamma_{0}}\in\big{(}{0,1/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ is a constant. Then, for any $T\in{\mathbb{N}_{+}}$,
@@ A. Useful Lemmas
Some preliminary results are given in this subsection. We first provide some results on the projection in the following lemma.
@@
In addition, let $\Phi(y)={\|{b-y}\|^{2}}+2\langle{a,y}\rangle$, then we know $\Phi$ is a strongly convex function with convexity parameter $\sigma=2$ and ${x_{c}}=\mathop{\arg\min}\limits_{y\in\mathbb{K}}\Phi(y)$. We can obtain
$\|{{x_{c}}-b}\|\leq\|a\|.$ (30)
@@
First, we know (29) holds from Lemma 3 in [35]. Then, Since $\Phi$ is a strongly convex function with convexity parameter $\sigma=2$, we have
$\Phi(b)\geq\Phi({{x_{c}}})+{\big{(}{\nabla\Phi({{x_{c}}})}\big{)}^{T}}({b-{x_{c}}})+\frac{\sigma}{2}\|{b-{x_{c}}}\|.$
From the optimality condition, we have
${\big{(}{\nabla\Phi({{x_{c}}})}\big{)}^{T}}({b-{x_{c}}})\geq 0.$
@@
$\Phi(b)\geq\Phi({{x_{c}}})+\frac{\sigma}{2}\|{b-{x_{c}}}\|.$
From $\Phi(y)={\|{b-y}\|^{2}}+2\langle{a,y}\rangle$, we have
$2\langle{a,b}\rangle\geq{\|{b-{x_{c}}}\|^{2}}+2\langle{a,{x_{c}}}\rangle+\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}.$
Thus, we have
$2\langle{a,b-{x_{c}}}\rangle\geq{\|{b-{x_{c}}}\|^{2}}+\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}.$
From the Cauchy-Schwarz inequality, we have
$2\langle{a,b-{x_{c}}}\rangle\leq 2\|a\|\|{b-{x_{c}}}\|.$
Thus, we have
${\|{b-{x_{c}}}\|^{2}}+\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}\leq 2\|a\|\|{b-{x_{c}}}\|.$
Due to $\sigma=2$, we have
$\|{b-{x_{c}}}\|\leq\|a\|.$
Therefore, we know that (30) holds. ∎
@@
From the Cauchy-Schwarz inequality and (31f), we have
@@
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]\leq{p{G_{1}}{\alpha_{t}}+p{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}.$ (47)
@@
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]$ $={\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{{z_{i,t+1}}-{x_{i,t}}}\|]\leq{{\alpha_{t}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{{{\hat{\omega}}_{i,t+1}}}\|]}$
$={{\alpha_{t}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\hat{\partial}{f_{i,t}}({{x_{i,t}}})+\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}}}\|]}$
$\leq{{\alpha_{t}}}\big{(}{p{G_{1}}+p{G_{2}}{\gamma_{t}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}$
$={p{G_{1}}{\alpha_{t}}+p{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}.$ (64)
where the second equality holds due to (15b); the last inequality holds due to (31c), (31f) and (51); the last equality holds due to ${\gamma_{t}}{\alpha_{t}}={\gamma_{0}}$. It follows form (64) that (47) holds. ∎
@@
${n_{T}}=\sum\limits_{t=1}^{T}{{\tilde{\varepsilon}_{2}^{2}{{{{G_{1}^{2}}}}}{\alpha_{t}}}}+{m_{T}},{\varepsilon_{2}}=\frac{{4\max\{{1,G_{2}^{2}{{\tilde{\varepsilon}}_{4}}}\}}}{{\min\{{1,\frac{1}{{2{\gamma_{0}}}}}\}}},$
${\varepsilon_{3}}={np{{G_{1}}}{G_{2}}{{\tilde{\varepsilon}}_{2}}},{\varepsilon_{4}}={pG_{2}^{2}{{\tilde{\varepsilon}}_{2}}{\gamma_{0}}+1}.$
@@
$\leq n{\tilde{\varepsilon}_{1}}+{{{\tilde{\varepsilon}}_{2}}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\big{(}{p{G_{1}}{\alpha_{t}}+p{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}}}.$ (76)